Circular No. 7849

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001jy, 2001jz, 2001ka, 2001kb, 2001kc, 2001kd B. Barris, University of Hawaii, on behalf of the High-Z Supernova

Search Team (cf. IAUC 7801 and references therein), reports the discovery of several apparent supernovae on I-band images taken in late 2001 (cf. IAUC 7802, 7805, 7806, and above citations), each of which has now faded from view. The host-galaxy emission for SN 2001jy and SN 2001kd yields z = 0.18 and 0.937, respectively. Each apparent supernova is within 2′′ of its host galaxy’s center. None of the apparent supernovae are present in images taken on 2002 Feb. 17 (limiting mag I = 25.8) with the Subaru telescope (+ SUPRIME) by K. Chambers and M. Liu. The I_{1} magnitudes below are for the discovery date; the I_{2} magnitudes are from observations made on 2001 Sept. 24, Oct. 9, 16, Nov. 12, 18, and 18 (respectively).

SN  2001 UT  α_{2000}  δ_{2000}  I_{1}  I_{2}
2001jy Sept. 12  4^{h}38^{m}54.^{s}67 − 1^{o}20′02′′.8 21.5 21.5
2001jz Sept. 24  4 39 20.82 − 1 37 48.0 23.9 23.3
2001ka Oct. 9  8 49 07.94 +44 10 00.2 21.6 22.0
2001kb Oct. 22  4 39 12.42 − 1 21 24.1 24.5 24.8
2001kc Nov. 12  7 50 01.97 +10 19 38.7 24.4 24.9
2001kd Nov. 12  7 50 31.24 +10 21 07.3 24.3 24.5

SUPERNOVA 2002bo IN NGC 3190 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum (range 370–750 nm) of SN 2002bo, obtained by V. Hradecky on Mar. 10.25 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a reddened and possibly peculiar type-Ia supernova well before maximum. Adopting the NED recession velocity of 1271 km/s for the host galaxy, the supernova expansion velocity is ∼ 17 900 km/s for Si II (rest 635.5 nm). This expansion velocity implies an age of ≈ 2 weeks before maximum, based upon comparison with values presented by Jha et al. (1999, Ap.J. Supp. 125, 73). The spectrum shows strong, narrow Na I D absorptions at the velocity of the host galaxy (0.3 nm EW). The Schlegel et al. (1998, Ap.J. 500, 525) dust maps indicate a Galactic reddening of E(B − V) = 0.025. Several titanium absorption features appear stronger than usual, indicating that this may be a subluminous event similar to SN 1999by (see Garnavich et al. 2002, Ap.J., in

press, http://xxx.lanl.gov/abs/astro-ph/?0105490).

2002 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green